# Logistic‑Gated Operators Enable Auditable Unit‑Aware Thresholds in Symbolic Regression

### v1

@@ Results
We first summarize predictive accuracy across methods and datasets to situate LGO within the current symbolic-regression landscape. Figure S1 visualizes the distribution of performance over up to ten random seeds per method and dataset (Hydraulic excludes two anomalous seeds (R2 for regression, AUROC for CTG); see SI Table S10). Each violin encodes the full seed distribution, black dots mark individual runs, and the numbers beneath each violin report the mean $\pm$ standard deviation across seeds.
@@
Figure S1: Performance comparison across methods and datasets. Violin plots show the distribution over up to ten seeds per method (Hydraulic excludes 2 anomalous seeds; see SI Table S10). The black dots are individual seeds; violin width reflects probability density. The baseline metric is R2 for regression datasets and AUROC for CTG; the mean and standard deviation for each violin are printed underneath (mean on the top line, $\pm$ std on the bottom line). The figure provides a high-level accuracy overview before parsimony and interpretability analyses.
@@
Across 10 random seeds, RILS-ROLS attains the highest accuracy, while LGOhard consistently outperforms LGOsoft and is competitive with strong baselines. Table S1 reports mean $\pm$ std on the held-out test set. Concretely, LGOhard reaches $R^{2}=0.800\pm 0.068$ with $\mathrm{RMSE}=0.817\pm 0.118$ and $\mathrm{MAE}=0.664\pm 0.104$, surpassing LGOsoft ($R^{2}=0.699\pm 0.129$). This confirms that explicit (hard) gating better captures the ICU’s switch-like dynamics.
Table S1: ICU composite risk score: mean $\pm$ std (10 seeds).
@@
RILS-ROLS again leads in $R^{2}$, followed by Operon, PySR and LGO variants (Table S2). Within LGO, LGOhard ($R^{2}=0.599\pm 0.176$) remains superior to LGOsoft ($0.549\pm 0.116$), though LGObase is strongest among LGO variants on this dataset ($0.677\pm 0.061$).
Table S2: NHANES metabolic score: mean $\pm$ std (10 seeds).
@@
LGO recovers thresholds in natural units post hoc. For ICU, the LGOhard lactate threshold concentrates around $1.89$ mmol/L (IQR [0.79, 2.07][0.79,\,2.07]). For NHANES, systolic_bp centers at $\sim 128.3$ mmHg and fasting glucose at $\sim 85.5$ mg/dL (Table S3). These values are consistent with the intuition that LGO discovers clinically plausible cutpoints from data.
Table S3: Representative thresholds from LGOhard (median and IQR in natural units).
@@
On ICU, LGO recovers credible gates for perfusion and hemodynamics: Lactate (mmol/L) and MAP (mmHg) align tightly with anchors, while Respiratory rate (min-1) lies in the yellow band, consistent with a more sensitive monitoring threshold. On NHANES, gates on SBP (mmHg), HDL (mg/dL) and Waist circumference (cm) closely match cardiometabolic anchors (green), and Fasting glucose (mg/dL) indicates an earlier, sensitivity-oriented turning point (yellow). Interquartile ranges (IQR) are narrow for SBP/HDL/Waist, indicating stability of the recovered gates across runs; see Figure S2 for the joint view of agreement and dispersion, and Table S4 for numerical values.
Beyond anchor agreement, we assess the stability of recovered thresholds across the ten random seeds. For each anchored feature we compute the seed-wise median threshold in natural units and summarize dispersion via interquartile ranges (IQR) and a bootstrap 95% CI. Across ICU and NHANES, SBP (mmHg), HDL (mg/dL), and waist circumference (cm) exhibit narrow IQRs, while fasting glucose (mg/dL) shows slightly wider variability yet remains centered in the yellow band relative to the diagnostic anchor — consistent with a conservative, screening-oriented turning point rather than a diagnostic cutoff. These stability views complement Figure S2 and Table S4 by showing not only where the gates land, but how consistently they are recovered across runs. These per‑seed medians in Tables S3 and S4 differ from the top‑k pool summaries in the SI (Tables S5 – S7), so identical medians/IQRs are not expected across the two views.
Figure S2: LGO threshold alignment with clinical guidelines (ICU and NHANES). Panels A and C: agreement heatmaps (green $\leq 10\%$, yellow $\leq 20\%$, red $20\%$). Only anchored features with valid median thresholds are shown, so grey (N/A) cells are suppressed. Each cell reports the median threshold (natural units) and the relative deviation $\Delta$. Panels B and D: distribution of the same thresholds – horizontal bars show the IQR, black dots mark medians, and red vertical lines mark anchors; the $x$-axis is in natural units. The figure conveys agreement (A,C) and stability (B,D) without duplicating the numerical details tabulated in Table S4.
Table S4: LGO-discovered thresholds (median [Q1, Q3]) versus domain anchors (ICU and NHANES). Values are in natural units.
@@
Figure S3: Gating mechanism usage across datasets. Top panels (one subplot per dataset): for LGOsoft and LGOhard, the blue bars show gate usage % (fraction of the top-$100$ models that include at least one LGO gate), and the orange bars show the median number of gates per model (zeros included). Bottom panel: gate usage % aggregated side-by-side across datasets, contrasting LGOsoft (light) vs. LGOhard (dark). Hard gates keep usage high where thresholding is needed (ICU, NHANES, CTG) while pruning superfluous gates (e.g., on Cleveland), and they realize fewer gates per model than soft gates, supporting parsimony and auditability.
@@
Figure S3 summarizes these measures across datasets and variants. Consistent with our interpretability hypothesis, LGOhard retains a high usage rate where thresholding is truly beneficial (e.g., ICU risk, NHANES metabolic) but prunes gates entirely when the signal does not support them (e.g., in a subset of Cleveland models). When gates are used, LGOhard typically realizes fewer gates per model than LGOsoft, yielding a sparser switching structure without sacrificing the ability to form crisp, auditable rules. This pattern aligns with our threshold audit (previous subsection): on ICU, gates for Lactate (mmol/L) and MAP (mmHg) tightly match clinical anchors (green), while Respiratory rate (min-1) lands in the yellow band ($27$ vs. $24$, $\Delta 12\%$), indicating a more sensitive monitor setting rather than a diagnostic cutoff.
Hard gates indeed discover sparser switching structure (Table S5): across datasets the median number of gates is consistently lower for LGOhard than for LGOsoft (e.g., ICU: $4.0$ vs. $10.0$; NHANES: $5.0$ vs. $12.5$). Combined with the threshold-alignment analysis, these results indicate that LGOhard uses gates selectively and economically — enough to capture clinically/physically meaningful cut-points, yet few enough to keep models compact and auditable.
Table S5: Gate usage by dataset: median number of gates among top-100 results per dataset.
@@
Beyond aggregate statistics, we inspect the top‑ranked model per method to visualize what is actually learned. Tables S6 and S7 juxtapose the best equation from LGOhard against strong baselines on ICU and NHANES. Two patterns emerge. First, where thresholds govern the phenomenon (NHANES cardiometabolic risk), LGOhard exposes explicit, unit‑aware gates on triglycerides, fasting glucose, SBP, HDL, and waist circumference, aligning with our threshold‑audit analysis. Second, when smooth trends suffice (ICU top‑1 in this split), LGOhard prunes gates and reduces to smooth primitives — consistent with our parsimony analysis — while baselines also remain smooth and non‑auditable. These exemplars complement the gate‑usage (Figure S3) and threshold‑alignment (Figure S2) results: gates are used selectively, and when used, they yield auditable cut‑points in natural units.
Table S6: ICU (composite risk): top‑1 equations by method (simplified forms for readability).
@@
Table S7: NHANES (metabolic score): top‑1 equations by method (simplified forms for readability).
@@
and audited in natural units (see Figure S2); $g(\cdot)$ indicates a smooth subexpression.
@@
To isolate LGO components, we compare three operator sets inside the same engine: base (no gates), LGOsoft, and LGOhard. Table S8 summarizes $R^{2}$ medians (IQR) over 10 seeds. LGOhard outperforms LGOsoft on ICU and Cleveland, and is closer to (or slightly above) base on threshold-heavy tasks; on Hydraulic, introducing hard gates can hurt when relations are predominantly smooth.
Table S8: Ablation within LGO: medians (IQR) of $R^{2}$ over 10 seeds.
@@ Discussion
Agreement with domain anchors (Figure S2; TableS4) substantiates this claim. On ICU and NHANES, $\approx 71\%$ (5/7) of assessed gates fall within $\leq 10\%$ and 100% within $\leq 20\%$ of guideline thresholds in natural units, with no red cells. The ICU gates for Lactate (mmol/L) and MAP (mmHg) align tightly with Sepsis/critical‑care anchors (2.0 mmol/L and 65 mmHg, respectively) 7, 8, while NHANES gates concur with cardiometabolic anchors for SBP, HDL and waist circumference 9, 10, 11. The fasting‑glucose gate appears earlier than the diagnostic cutoff (yellow band), surfacing a conservative, sensitivity‑oriented turning point — useful for screening workflows.
Parsimony of gating (Figure S3; TableS5) provides a second line of evidence. Across datasets, LGOhard uses fewer gates per model than LGOsoft (e.g., ICU 4.0 vs 10.0; NHANES 5.0 vs 12.5 median gates among top‑100 models), while maintaining high usage rates only where thresholding is supported by data. This selective economy is consistent with our ablations (Table S8): hard gates outperform or match soft gates on threshold‑heavy problems (ICU, Cleveland), and are near base on smoother tasks.
A third perspective is competitive accuracy with distributional stability (Figure S1). Although our goal is not to chase SOTA point estimates, LGO sits within the competitive envelope across datasets relative to strong SR baselines (RILS-ROLS, Operon, PySR). The violins reveal reasonable dispersion across seeds, and when a small loss in pure accuracy occurs, it is compensated by the gains in auditability and sparsity — a trade‑off explicitly desired in scientific and clinical modeling 3.
@@
Yellow-band deviations (e.g., fasting glucose in NHANES; Figure S2C) are best read as earlier change-points rather than errors. For risk stratification and routine monitoring, a conservative turning point can be preferable to a diagnostic cutoff: the former flags emerging risk, while the latter certifies disease. We therefore explicitly separate screening-appropriate gates — often earlier and yellow — from diagnostic thresholds—typically green—so that model interpretation follows clinical workflow.
@@
All code to reproduce results (LGO primitives; typed GP engine; baselines PySR, Operon, PSTree, RILS–ROLS; plotting and aggregation scripts) is available on GitHub: https://github.com/oudeng/LGO. A camera-ready snapshot will be archived on Zenodo ([ZENODO_DOI_URL]). Reproducibility details (software environments, artifact index, operator sets/budgets) are summarized in SI §1; Tables S1–S3.
@@ Acknowledgments
The work was supported in part by the 2022-2024 Masaru Ibuka Foundation Research Project on Oriental Medicine, 2020-2025 JSPS A3 Foresight Program (Grant No. JPJSA3F20200001), 2022–2024 Japan National Initiative Promotion Grant for Digital Rural City, 2023 and 2024 Waseda University Grants for Special Research Projects (Nos. 2023C-216 and 2024C-223), 2023-2024 Waseda University Advanced Research Center Project for Regional Cooperation Support, and 2023-2024 Japan Association for the Advancement of Medical Equipment (JAAME) Grant.

### v2

@@ Results
We first summarize predictive accuracy across methods and datasets to situate LGO within the current symbolic-regression landscape. Figure 1 visualizes the distribution of performance over up to ten random seeds per method and dataset (Hydraulic excludes two anomalous seeds (R2 for regression, AUROC for CTG); see SI Table S10). Each violin encodes the full seed distribution, black dots mark individual runs, and the numbers beneath each violin report the mean $\pm$ standard deviation across seeds.
@@
Figure 1: Performance comparison across methods and datasets. Violin plots show the distribution over up to ten seeds per method (Hydraulic excludes 2 anomalous seeds; see SI Table S10). The black dots are individual seeds; violin width reflects probability density. The baseline metric is R2 for regression datasets and AUROC for CTG; the mean and standard deviation for each violin are printed underneath (mean on the top line, $\pm$ std on the bottom line). The figure provides a high-level accuracy overview before parsimony and interpretability analyses.
@@
Across 10 random seeds, RILS-ROLS attains the highest accuracy, while LGOhard consistently outperforms LGOsoft and is competitive with strong baselines. Table 1 reports mean $\pm$ std on the held-out test set. Concretely, LGOhard reaches $R^{2}=0.800\pm 0.068$ with $\mathrm{RMSE}=0.817\pm 0.118$ and $\mathrm{MAE}=0.664\pm 0.104$, surpassing LGOsoft ($R^{2}=0.699\pm 0.129$). This confirms that explicit (hard) gating better captures the ICU’s switch-like dynamics.
Table 1: ICU composite risk score: mean $\pm$ std (10 seeds).
@@
RILS-ROLS again leads in $R^{2}$, followed by Operon, PySR and LGO variants (Table 2). Within LGO, LGOhard ($R^{2}=0.599\pm 0.176$) remains superior to LGOsoft ($0.549\pm 0.116$), though LGObase is strongest among LGO variants on this dataset ($0.677\pm 0.061$).
Table 2: NHANES metabolic score: mean $\pm$ std (10 seeds).
@@
LGO recovers thresholds in natural units post hoc. For ICU, the LGOhard lactate threshold concentrates around $1.89$ mmol/L (IQR [0.79, 2.07][0.79,\,2.07]). For NHANES, systolic_bp centers at $\sim 128.3$ mmHg and fasting glucose at $\sim 85.5$ mg/dL (Table 3). These values are consistent with the intuition that LGO discovers clinically plausible cutpoints from data.
Table 3: Representative thresholds from LGOhard (median and IQR in natural units).
@@
On ICU, LGO recovers credible gates for perfusion and hemodynamics: Lactate (mmol/L) and MAP (mmHg) align tightly with anchors, while Respiratory rate (min-1) lies in the yellow band, consistent with a more sensitive monitoring threshold. On NHANES, gates on SBP (mmHg), HDL (mg/dL) and Waist circumference (cm) closely match cardiometabolic anchors (green), and Fasting glucose (mg/dL) indicates an earlier, sensitivity-oriented turning point (yellow). Interquartile ranges (IQR) are narrow for SBP/HDL/Waist, indicating stability of the recovered gates across runs; see Figure 2 for the joint view of agreement and dispersion, and Table 4 for numerical values.
Beyond anchor agreement, we assess the stability of recovered thresholds across the ten random seeds. For each anchored feature we compute the seed-wise median threshold in natural units and summarize dispersion via interquartile ranges (IQR) and a bootstrap 95% CI. Across ICU and NHANES, SBP (mmHg), HDL (mg/dL), and waist circumference (cm) exhibit narrow IQRs, while fasting glucose (mg/dL) shows slightly wider variability yet remains centered in the yellow band relative to the diagnostic anchor — consistent with a conservative, screening-oriented turning point rather than a diagnostic cutoff. These stability views complement Figure 2 and Table 4 by showing not only where the gates land, but how consistently they are recovered across runs. These per‑seed medians in Tables 3 and 4 differ from the top‑k pool summaries in the SI (Tables S5 – S7), so identical medians/IQRs are not expected across the two views.
Figure 2: LGO threshold alignment with clinical guidelines (ICU and NHANES). Panels A and C: agreement heatmaps (green $\leq 10\%$, yellow $\leq 20\%$, red $20\%$). Only anchored features with valid median thresholds are shown, so grey (N/A) cells are suppressed. Each cell reports the median threshold (natural units) and the relative deviation $\Delta$. Panels B and D: distribution of the same thresholds – horizontal bars show the IQR, black dots mark medians, and red vertical lines mark anchors; the $x$-axis is in natural units. The figure conveys agreement (A,C) and stability (B,D) without duplicating the numerical details tabulated in Table 4.
Table 4: LGO-discovered thresholds (median [Q1, Q3]) versus domain anchors (ICU and NHANES). Values are in natural units.
@@
Figure 3: Gating mechanism usage across datasets. Top panels (one subplot per dataset): for LGOsoft and LGOhard, the blue bars show gate usage % (fraction of the top-$100$ models that include at least one LGO gate), and the orange bars show the median number of gates per model (zeros included). Bottom panel: gate usage % aggregated side-by-side across datasets, contrasting LGOsoft (light) vs. LGOhard (dark). Hard gates keep usage high where thresholding is needed (ICU, NHANES, CTG) while pruning superfluous gates (e.g., on Cleveland), and they realize fewer gates per model than soft gates, supporting parsimony and auditability.
@@
Figure 3 summarizes these measures across datasets and variants. Consistent with our interpretability hypothesis, LGOhard retains a high usage rate where thresholding is truly beneficial (e.g., ICU risk, NHANES metabolic) but prunes gates entirely when the signal does not support them (e.g., in a subset of Cleveland models). When gates are used, LGOhard typically realizes fewer gates per model than LGOsoft, yielding a sparser switching structure without sacrificing the ability to form crisp, auditable rules. This pattern aligns with our threshold audit (previous subsection): on ICU, gates for Lactate (mmol/L) and MAP (mmHg) tightly match clinical anchors (green), while Respiratory rate (min-1) lands in the yellow band ($27$ vs. $24$, $\Delta 12\%$), indicating a more sensitive monitor setting rather than a diagnostic cutoff.
Hard gates indeed discover sparser switching structure (Table 5): across datasets the median number of gates is consistently lower for LGOhard than for LGOsoft (e.g., ICU: $4.0$ vs. $10.0$; NHANES: $5.0$ vs. $12.5$). Combined with the threshold-alignment analysis, these results indicate that LGOhard uses gates selectively and economically — enough to capture clinically/physically meaningful cut-points, yet few enough to keep models compact and auditable.
Table 5: Gate usage by dataset: median number of gates among top-100 results per dataset.
@@
Beyond aggregate statistics, we inspect the top‑ranked model per method to visualize what is actually learned. Tables 6 and 7 juxtapose the best equation from LGOhard against strong baselines on ICU and NHANES. Two patterns emerge. First, where thresholds govern the phenomenon (NHANES cardiometabolic risk), LGOhard exposes explicit, unit‑aware gates on triglycerides, fasting glucose, SBP, HDL, and waist circumference, aligning with our threshold‑audit analysis. Second, when smooth trends suffice (ICU top‑1 in this split), LGOhard prunes gates and reduces to smooth primitives — consistent with our parsimony analysis — while baselines also remain smooth and non‑auditable. These exemplars complement the gate‑usage (Figure 3) and threshold‑alignment (Figure 2) results: gates are used selectively, and when used, they yield auditable cut‑points in natural units.
Table 6: ICU (composite risk): top‑1 equations by method (simplified forms for readability).
@@
Table 7: NHANES (metabolic score): top‑1 equations by method (simplified forms for readability).
@@
and audited in natural units (see Figure 2); $g(\cdot)$ indicates a smooth subexpression.
@@
To isolate LGO components, we compare three operator sets inside the same engine: base (no gates), LGOsoft, and LGOhard. Table 8 summarizes $R^{2}$ medians (IQR) over 10 seeds. LGOhard outperforms LGOsoft on ICU and Cleveland, and is closer to (or slightly above) base on threshold-heavy tasks; on Hydraulic, introducing hard gates can hurt when relations are predominantly smooth.
Table 8: Ablation within LGO: medians (IQR) of $R^{2}$ over 10 seeds.
@@ Discussion
Agreement with domain anchors (Figure 2; Table4) substantiates this claim. On ICU and NHANES, $\approx 71\%$ (5/7) of assessed gates fall within $\leq 10\%$ and 100% within $\leq 20\%$ of guideline thresholds in natural units, with no red cells. The ICU gates for Lactate (mmol/L) and MAP (mmHg) align tightly with Sepsis/critical‑care anchors (2.0 mmol/L and 65 mmHg, respectively) 7, 8, while NHANES gates concur with cardiometabolic anchors for SBP, HDL and waist circumference 9, 10, 11. The fasting‑glucose gate appears earlier than the diagnostic cutoff (yellow band), surfacing a conservative, sensitivity‑oriented turning point — useful for screening workflows.
Parsimony of gating (Figure 3; Table5) provides a second line of evidence. Across datasets, LGOhard uses fewer gates per model than LGOsoft (e.g., ICU 4.0 vs 10.0; NHANES 5.0 vs 12.5 median gates among top‑100 models), while maintaining high usage rates only where thresholding is supported by data. This selective economy is consistent with our ablations (Table 8): hard gates outperform or match soft gates on threshold‑heavy problems (ICU, Cleveland), and are near base on smoother tasks.
A third perspective is competitive accuracy with distributional stability (Figure 1). Although our goal is not to chase SOTA point estimates, LGO sits within the competitive envelope across datasets relative to strong SR baselines (RILS-ROLS, Operon, PySR). The violins reveal reasonable dispersion across seeds, and when a small loss in pure accuracy occurs, it is compensated by the gains in auditability and sparsity — a trade‑off explicitly desired in scientific and clinical modeling 3.
@@
Yellow-band deviations (e.g., fasting glucose in NHANES; Figure 2C) are best read as earlier change-points rather than errors. For risk stratification and routine monitoring, a conservative turning point can be preferable to a diagnostic cutoff: the former flags emerging risk, while the latter certifies disease. We therefore explicitly separate screening-appropriate gates — often earlier and yellow — from diagnostic thresholds—typically green—so that model interpretation follows clinical workflow.
@@
All code to reproduce results (LGO primitives; typed GP engine; baselines PySR, Operon, PSTree, RILS–ROLS; plotting and aggregation scripts) is available on GitHub: https://github.com/oudeng/LGO. A camera-ready snapshot will be archived on Zenodo (DOI:10.5281/zenodo.17299423). Reproducibility details (software environments, artifact index, operator sets/budgets) are summarized in SI §1; Tables S1–S3.
@@ Acknowledgments
The work was supported in part by the 2022-2024 Masaru Ibuka Foundation Research Project on Oriental Medicine, 2020-2025 JSPS A3 Foresight Program (Grant No. JPJSA3F20200001), 2022-2024 Japan National Initiative Promotion Grant for Digital Rural City, 2023 and 2024 Waseda University Grants for Special Research Projects (Nos. 2023C-216 and 2024C-223), 2023-2024 Waseda University Advanced Research Center Project for Regional Cooperation Support, and 2023-2024 Japan Association for the Advancement of Medical Equipment (JAAME) Grant.